Find the equation of the tangent line at the given value of x on the curve.
y^3+xy^2-25=x+2y^2 ; x=2
1
Expert's answer
2020-04-13T19:33:37-0400

Let's find "\frac{dy}{dx}"

"3y^2y'+y^2+2xyy'=1+4yy'"

"y'(3y^2+2xy-4y)=1-y^2"

"y'=\frac{1-y^2}{y(3y+2x-4)}"

Let's find

"y(2)=y_0:y_0^3+2y_0^2-25=2+2y_0^2"

"y_0^3=27;y_0=3"

"y'(2)=\frac{1-3^2}{3(3*3+2*2-4)}=-\frac{8}{27}"

The equation of the tangent line

"y=y'(2)(x-2)+y_0"

Therefore

"y_t=-\frac{8}{27}(x-2)+3"